OpenStudy (anonymous):

A laptop computer purchased for $1400. After each year the resale value decreased by 35%. What will be the resale value after 4 years?

OpenStudy (anonymous):

1400(1-.35)^4

OpenStudy (anonymous):

Year 1 (35% of 1400= 490) 1400- 490= 910 Year 2 (35% of 910 = 318.5) 910- 318.5= 591.5 Year 3 (35% of 591.5= 207 * rounded to the nearest hundredth*) 591.5- 207= 384.5 Year 4 (35% of 384.5 = 134.6 *rounded to the nearest hundredth* ) 384.5 - 134.6= 249.9.

OpenStudy (anonymous):

@flapdragon why do such long calculation you will get same answer as mine unless you haven't made any mistake

OpenStudy (anonymous):

Just in case the person who asked the question needs to know how I got the answer. :)

OpenStudy (anonymous):

in this case a=1,b=k,c=k+1 so for \[\Delta = k ^{2}-4\left( k +1 \right) =k ^{2}-4k -4\] now use quadratic formula on this to find k.\[\div 2k=\left( 4\pm \sqrt{16-4\times -4} \right)\div2=\left( 4\pm \sqrt{32} \right)\div2=\left( 4\pm4\sqrt{2} \right)\]\[\div2=2\pm2\]